OpenStudy (studybird123):

Solve for the variable. z/5 = 6/30 A. 1/5 B. 1 C. 5 D. 25

OpenStudy (studybird123):

Is it B.

OpenStudy (anonymous):

z = 6/30 * 5 z = 30/30 z = 1 so it's B